OpenStudy (anonymous):

PLEASE HELP! The length of a rectangular playing field is 7 yards longer than triple the width. If the perimeter of the rectangular playing field is 438 yards, what are it's dimensions in yards?

OpenStudy (anonymous):

\[width=x\]\[length=3x+7\]\[438=2(x)+2(3x+7)\]\[438=2x+6x+14\]\[424=8x\]\[53=x\]\[width=53yards\]\[length=3(53)+7=166yards\]